A cubical stero speaker measures 35 centimeters along each edge. The expression for finding the surface area of a cube is 6s squared where s is the length of each edge. Find the surface area of the stero speaker.
Answer by stanbon(75887)

A cubical stero speaker measures 35 centimeters along each edge. The expression for finding the surface area of a cube is 6s squared where s is the length of each edge. Find the surface area of the stero speaker.
----
SA = 6*35^2 = 7.35x10^3 = 7350 sq cm
=================
Cheers,
Stan H.
=================